OpenStudy (anonymous):

1/6-2/3m> 1/4 1/6-2/3m> 1/4 @Mathematics

OpenStudy (agreene):

-8<m<0

OpenStudy (anonymous):

multiply everything by 3m 3m/6-6m/3m>3m/4 m/2-2>3m/4 multiple everything by 4 2m-8>3m -8>m

OpenStudy (anonymous):

\[\frac{1}{6}-\frac{2}{3}m>\frac{1}{4}\] multiply by 12 to get \[2-8m>3\] add 8m get \[2>8m+3\] subtract 3 get \[-1>8m\] divide by 8 get \[-\frac{1}{8}>m\]

OpenStudy (anonymous):

really getting some random answers this evening

OpenStudy (anonymous):

i don't see where i went wrong

OpenStudy (anonymous):

@satellite73 I know right :P @da1bearsfan I don't even know what I am doing..

OpenStudy (anonymous):

@satellite73 how did u get multiply by 12 though?

OpenStudy (agreene):

is the question: \[\frac{1}{6}-\frac{2}{3m}>\frac{1}{4}\] OR \[\frac{1}{6}-\frac{2}{3}m>\frac{1}{4}\] in the first case, my answer is correct, and in the latter satellite is.

OpenStudy (anonymous):

you cannot multiply by 3m because you don't know whether it is positive or negative

OpenStudy (anonymous):

@ nico i multiplied by 12 so i didn't have to work with the fractions. you don't need to do that if you don't want

OpenStudy (anonymous):

Ohh ok now i know my teacher never explained that to me

OpenStudy (anonymous):

btw it really was \[\frac{1}{6}-\frac{2}{3}m>\frac{1}{4}\] right?

OpenStudy (anonymous):

yes it was satellite

OpenStudy (anonymous):

you can subtract \[\frac{1}{6} \] if you want and get \[-\frac{2}{3}m>\frac{1}{4}-\frac{1}{6}\] \[-\frac{2}{3}m>\frac{1}{12}\] \[m<\frac{1}{12}\times -\frac{3}{2}\] \[m<-\frac{1}{8}\]

OpenStudy (anonymous):

when you multiply both sides by a negative number you have to switch the inequality. i find it easier to work without the fractions, that is why i multiplied EVERYTHING by 12